# On the numerical stability of sketched GMRES

### v1

@@ 5 Adaptively restarting sGMRES
Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S^{(i)}\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$. Then there exists $k^{(i)}\leq n$ such that
$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl{(}\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr{)}\geq 1.$
@@
$\Lambda^{(i)}_{1}=O(\bm{u})\frac{1+\epsilon}{1-\epsilon}\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\kappa_{\mathsf{F}}(A)\ll 1$
holds for $i$. Then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
@@
As a consequence of Theorem 2, it is clear that a restarted variant of sketched GMRES (Algorithm 2) may allow us to recover the stability of sGMRES when the conditioning of the basis grows large.
It is clear from our analysis that care should be taken to ensure $\kappa(Z_{1:i})$ does not get too large as the iteration progresses. In fact, as noted in [10], due to the subspace embedding property (5), the condition number of $Z_{1:i}$ can be monitored inexpensively by explicitly computing the condition number of $SZ_{1:i}$. The numerical stability of sGMRES can thus be controlled by monitoring the condition number of $SZ_{1:i}$ in every iteration, and restarting when it grows above some specified tolerance $\texttt{tol}_{B}$. For the case of a truncated Arnoldi implementation, or the sketch-and-select Arnoldi implementations, the truncation parameter $t$ can then be increased upon restart to ensure the condition number of the basis $Z_{1:i}$ does not grow as rapidly as it did in the previous cycle.
As an alternative, due to the result of Lemma 2, we may also opt to monitor
$\tau_{i}:=\frac{\|Z_{1:i}\rVert\|y^{(i)}\rVert}{\|x^{(i)}\rVert},$
from (43) at each iteration, since it provides a sharper bound on the backward error, and can be inexpensively estimated as
$\tilde{\tau}_{i}:=\frac{\|SZ_{1:i}\rVert\|y^{(i)}\rVert}{\|Sx^{(i)}\rVert}.$
We may then opt to restart when $\texttt{tol}_{\tau}\cdot\tau_{i}\geq 1$, for some chosen tolerance $\texttt{tol}_{\tau}$. Again, optionally, we may then wish to increase the value of the truncation parameter $t$. In Algorithm 3, we present a hybrid approach, where we monitor $\tilde{\tau}_{i}$, and if it is large, we then restart. Before restarting, we also check if $SB_{1:i}$ is large, and if it is, we increase the truncation parameter, by, e.g., doubling it.
@@
1:  $r^{(0)}\leftarrow b-Ax^{(0)}$ and $g\leftarrow SM_{L}^{-1}r^{(0)}$.
2:  $\tilde{x}^{(0)}\leftarrow Sx^{(0)}$.
3:  $B_{1}\leftarrow r^{(0)}/\|r^{(0)}\rVert$.
@@
10:     $\tilde{Z}_{i}=SZ_{i}$ and $\tilde{x}^{(i)}\leftarrow\tilde{x}^{(0)}+\tilde{Z}_{1:i}y^{(i)}$.
@@
14:     $\tilde{\tau}_{i}\leftarrow\|\tilde{Z}_{1:i}\rVert\|y^{(i)}\rVert/\|\tilde{x}^{(i)}\rVert$
@@
16:        if $\texttt{tol}_{B}\cdot\kappa(\tilde{Z}_{1:i})\geq 1$ then
17:           $t\leftarrow 2t$
@@ 6 Numerical Experiments
In this section, we present numerical experiments that demonstrate the results of our stability analysis. All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM. We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [7]. Our code is available at https://github.com/burkel8/stability_sGMRES. The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [2]).
We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-saa, respectively, in the plots). Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [7], we use the sGMRES-saa-pinv method, as it was shown to produce the best conditioned basis in their experiments. We also ran the experiments with sGMRES-saa-greedy and obtained similar results, and thus do not include this method here. We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [7] for more details on these methods. In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method. All comparisons to standard GMRES use GMRES with modified Gram-Schmidt. We do not use any preconditioner in the experiments, i.e., $M_{L}=M_{R}=I$. Finally, we drop the $\hat{.}$ notation in this section, as all quantities are computed quantities.
@@ 6.1 Illustration of the backward error bound
In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5. For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [10]. We display the backward error, $\tau_{i}$ from (43), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration. These condition number approximations are denoted as $\kappa(B_{1:i})$ and $\kappa(AB_{1:i})$ in the plots.
@@ 6.1.1 Attainable worst case for sGMRES
The results are shown in Figure 1, where we see that the condition number of $B_{1:i}$ and $\tau_{i}$ generated by the truncated Arnoldi grows to $10^{16}$, leading to a poor backward error result for sGMRES-trunc. In the case of sGMRES-saa, the backward error is approximately $10^{-7}$ as both the condition number of $B_{1:i}$ and $\tau_{i}$ are not too large. For GMRES, it can be seen that the condition number of $B_{1:i}$ remains small, as does $\tau_{i}$, yielding a very small backward error after $400$ iterations. Thus, we can see that the bound in Theorem 1 can be attainable for some cases.
@@ 6.1.2 Comparison between the bounds shown in Theorem 1 and Lemma 2
As shown in Figure 2, the behavior of the backward error for both sGMRES-trunc and sGMRES-saa diverges; notably, while the backward error for sGMRES-saa stops improving after about 160 iterations, sGMRES-trunc continues to reduce its error reaching a small level, despite both methods having ill-conditioned $B_{1:i}$. This indicates that for this scenario, the bounds shown by Theorem 1, specifically (29) and (39), are overly pessimistic and fail to account for the discrepancy observed. We further notice that $\tau_{i}$ for sGMRES-trunc is considerably smaller than for sGMRES-saa, explaining the superior backward error performance of sGMRES-trunc, as anticipated by Lemma 2.
@@ 6.1.3 Comparison between different sGMRES variants with and without restarting
In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [4] (also tested in [7]). The right-hand side vector is randomly generated according to the standard normal distribution. We run $\texttt{nrestarts}=5$ cycles of restarted sGMRES with the truncated Arnoldi process, and sketch-and-select Arnoldi, with $m=150$ iterations per cycle, totaling $750$ iterations. We also run restarted GMRES with the same parameters. Additionally, we perform sGMRES with truncated Arnoldi and sketch-and-select Arnoldi with $m=750$ iterations (without restarts). All implementations of sGMRES use a truncation parameter $t=3$.
@@
6.2 sGMRES with the adaptively restarting strategy
In the final experiment, we test the various adaptive restarting strategies discussed in Section 3. In particular, we test
  • sGMRES (monitor $\tau_{i}$): monitor $\tau_{i}$, and when it gets too large, restart.
  • sGMRES (monitor $\kappa(B_{1:i})$): monitor the condition number of $B_{1:i}$, and when it gets too large, restart.
  • sGMRES (adapt monitor $\tau_{i}$): monitor $\tau_{i}$, and when it gets too large, double the truncation parameter $t$, and restart. Also restart and double $t$ every $m$ iterations even when $\tau_{i}$ is small.
  • sGMRES (adapt monitor $\kappa(B_{1:i})$): monitor the condition number of $B_{1:i}$, and when it gets too large, double the truncation parameter $t$, and restart. Also restart and double $t$ every $m$ iterations even when the condition number of $B_{1:i}$ is small.
  • sGMRES (adapt monitor $\tau_{i}$ and $\kappa(B_{1:i})$): monitor $\tau_{i}$, and when it gets too large, check the size of $\kappa(B_{1:i})$. If $\kappa(B_{1:i})$ is large, double the truncation parameter $t$, and restart. If $\kappa(B_{1:i})$ is not large enough, just restart without increasing $t$ (this method is shown in Algorithm 3). Also restart and double $t$ every $m$ iterations even when the condition number of $B_{1:i}$ is small.
@@
We solve a linear system by restarted sGMRES with the truncated Arnoldi process using all the various restarting strategies (beginning with $t=2$ for each method). We then plot the backward error curves produced by sGMRES with each restarting strategy, alongside the backward error curves produced by standard restarted GMRES. In each example, the right-hand side is randomly generated according to the standard normal distribution, and both $\texttt{tol}_{\tau}$ and $\texttt{tol}_{B}$ are taken to be the MATLAB machine precision obtained by the eps command.
For GMRES, we restart a maximum of $d$ times, every $m$ iterations. We then allow a maximum of $md$ iterations for all implementations of sGMRES methods, and each implementation to restart according to its own strategy.
In the first example, the matrix $A$ is the Norris/stomach matrix, $m=150$ and $d=4$. In the second example, the matrix $A$ is the fs_760_1 matrix obtained from SuiteSparse, and $m=30$ and $d=20$. In the final example, the matrix $A$ is taken to be the sherman2 matrix from SuiteSparse, with $m=200$ and $d=5$. The results are shown in Figure 4. The curve for sGMRES (monitor $\tau_{i}$) is hidden behind the curve for sGMRES (adapt monitor $\tau_{i}$) for each of the examples we tested. This is because $\tau_{i}$ is not directly dependent on the condition number of $B_{1:i}$, and so increasing $t$ does not help in this case. From the examples we have tested, sGMRES (adapt monitor $\tau_{i}$ and $\kappa(B_{1:i})$) seems to be the best approach as it produces backward error curves close to standard restarted GMRES in each of the examples.
Figure 4: Adaptive restarting experiment. From left to right, the backward error plots for Norris/Stomach, fs_760_1, and sherman2. The curve for sGMRES (monitor $\tau_{i}$), i.e., the dark blue line, is hidden behind the curve for sGMRES (adapt monitor $\tau_{i}$), i.e., the yellow line, for each of the examples we tested. The curve for sGMRES (adpt monitor $\tau_{i}$ and $\kappa(B_{1:i})$), i.e., the green line, is almost hidden behind the curve for GMRES, i.e., the light blue one.

### v2

@@ 5 Adaptive restarting sGMRES
Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$. Then there exists $k^{(i)}\leq n$ such that
$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl(\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr)\geq 1.$ (45)
@@
$\Lambda^{(i)}_{1}=\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\ll 1$ (46)
holds for $i$ with
$\bar{\tau}_{k^{(i)}}=\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}.$
Then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
@@
Proof.
Similar to the proof of [BHMV2024, Theorem 4.1], we will employ [BHMV2024, Lemma 4.2], which requires verifying the assumption [BHMV2024, (4.14)], i.e., bounding $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ by $\|b-A\hat{x}^{(i)}\rVert$ and $\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert$. From Lemma 2, we have
$\begin{split}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert&\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\biggl(\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}\|A\hat{d}^{(i)}\rVert+\|\hat{r}^{(i)}\rVert\biggr)\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\biggl(\bar{\tau}_{k^{(i)}}\bigl(\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert+\|\hat{r}^{(i)}\rVert\bigr)+\|\hat{r}^{(i)}\rVert\biggr)\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\biggl(\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\|\hat{r}^{(i)}\rVert+\bar{\tau}_{k^{(i)}}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert\biggr).\end{split}$ (48)
By moving the term related to $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ from the right-hand side of (48) to the left-hand side and using assumption (46), we further obtain
$\begin{split}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert&\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\|\hat{r}^{(i)}\rVert\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\bigl(\|b-A\hat{x}^{(i)}\rVert\\ &\quad+O(\bm{u})(\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert)\bigr),\end{split}$ (49)
which gives [BHMV2024, (4.14)]. Together with [BHMV2024, Lemma 4.2], we can draw the conclusion. ∎
Remark 6.
When no preconditioning is used, i.e., $M_{L}=M_{R}=I$ and $\hat{Z}_{1:k^{(i)}}^{(i)}=\hat{B}_{1:k^{(i)}}^{(i)}$, if it holds for the key dimension $k^{(i)}$ that
$\Lambda^{(i)}_{1}=O(\bm{u})\frac{1+\epsilon}{1-\epsilon}\biggl(1+\frac{\|\hat{B}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}\biggr)\ll 1,$ (50)
then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
Remark 7.
As a consequence of Theorem 2, it is clear that a restarted variant of sketched GMRES (Algorithm 2) may allow us to recover the stability of sGMRES when $\bar{\tau}_{k^{(i)}}$ is moderate. However, it should be noted that the assumptions, particularly (45), of this theorem are overly pessimistic. In general, selecting a sufficiently large $m$ is required to guarantee (45), but, in practice, we typically select a relatively small value for the maximum iterations per restart cycle, which makes it difficult to satisfy (45).
As discussed in Remark 7, for restarted sGMRES, there are two possible reasons for instability: One is due to choosing an $m$ that is too small; the other is due to using a value of $t$ that is too small, which has an influence on $\bar{\tau}_{k^{(i)}}$. Since the choice of $m$ depends mainly on storage in practice, here we only consider recovering from the instability caused by a small $t$. Employing a small $t$ can result in considerable acceleration, though it might also introduce instability issues.
Next, our aim is to propose an adaptive restarting strategy to recover from the instability of sGMRES caused by using a small $t$. The definition of $\Lambda_{1}^{(1)}$ in Theorem 2 relying on $\bar{\tau}_{k^{(i)}}$, which influences the reduced factor of the backward error in the $i$-th restart cycle, implies that care should be taken to ensure $\bar{\tau}_{k^{(i)}}$ does not get too large as the iteration progresses. In fact, as noted in [NT2024], due to the subspace embedding property (5), $\bar{\tau}_{i}$ can be inexpensively estimated as
$\tilde{\tau}_{i}:=\frac{\|SZ_{1:i}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|y^{(i)}\rVert}{\|SAZ_{1:i}y^{(i)}\rVert}.$
The numerical stability of sGMRES can thus be controlled by monitoring $\tilde{\tau}_{i}$ at each iteration, and doubling the value of $t$ when it grows above some specified tolerance. We then proceed to double $t$ again only if $\tilde{\tau}_{i}$ is still large and $\tilde{\tau}_{i}$ significantly exceeds $\tilde{\tau}_{i-1}$, indicating that $t$ remains insufficiently large. Restarted sGMRES combined with this adaptive strategy is presented in Algorithm 4.
@@
[…]
@@
10:  $\tilde{Z}_{i}=SZ_{i}$ and $\tilde{d}^{(i)}=C_{1:i}y^{(i)}$.
@@
14:  $\tilde{\tau}_{i}\leftarrow\|\tilde{Z}_{1:i}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|y^{(i)}\rVert/\|\tilde{d}^{(i)}\rVert$.
@@
16:   if $\tilde{\tau}_{i}>1.1\cdot\tilde{\tau}_{i-1}$ then
17:    $t\leftarrow\min\{i+1,2t\}$.
@@ 6 Numerical Experiments
In this section, we present numerical experiments that demonstrate the results of our stability analysis. All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM. We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [guttel2024sketch]. Our code is available at https://github.com/burkel8/stability_sGMRES. The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [balabanov2025randomized]).
We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-ssa, respectively, in the plots). Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [guttel2024sketch], we use the sGMRES-ssa-pinv method, as it was shown to produce the best conditioned basis in their experiments. We also ran the experiments with sGMRES-ssa-greedy and obtained similar results, and thus do not include this method here. We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [guttel2024sketch] for more details on these methods. In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method. All comparisons to standard GMRES use GMRES with modified Gram-Schmidt. We do not use any preconditioner in the experiments, i.e., $M_{L}=M_{R}=I$. Finally, we drop the $\hat{.}$ notation in this section, as all quantities are computed quantities.
@@ 6.1 Illustration of the backward error bound
In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5. For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [NT2024]. We display the backward error, $\tau_{i}$ from (44), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration. These condition number approximations are denoted as $\kappa(B_{1:i})$ and $\kappa(AB_{1:i})$ in the plots.
@@ 6.1.1 Attainable worst case for sGMRES
The results are shown in Figure 1, where we see that the condition number of $B_{1:i}$ and $\tau_{i}$ generated by the truncated Arnoldi grows to $10^{16}$, leading to a poor backward error result for sGMRES-trunc. In the case of sGMRES-ssa, the backward error is approximately $10^{-7}$ as both the condition number of $B_{1:i}$ and $\tau_{i}$ are not too large. For GMRES, it can be seen that the condition number of $B_{1:i}$ remains small, as does $\tau_{i}$, yielding a very small backward error after $400$ iterations. Thus, we can see that the bound in Theorem 1 can be attainable for some cases.
@@ 6.1.2 Comparison between the bounds shown in Theorem 1 and Lemma 2
As shown in Figure 2, the behavior of the backward error for both sGMRES-trunc and sGMRES-ssa diverges; notably, while the backward error for sGMRES-ssa stops improving after about 160 iterations, sGMRES-trunc continues to reduce its error reaching a small level, despite both methods having ill-conditioned $B_{1:i}$. This indicates that for this scenario, the bounds shown by Theorem 1, specifically (29) and (51), are overly pessimistic and fail to account for the discrepancy observed. We further notice that $\tau_{i}$ for sGMRES-trunc is considerably smaller than for sGMRES-ssa, explaining the superior backward error performance of sGMRES-trunc, as anticipated by Lemma 2. We also observe but do not find a clear theoretical explanation for the phenomenon that $\tau_{i}$ can even decrease after $\kappa(AB)$ and $\kappa(B)$ are both very ill-conditioned, which means that solving the least squares problem is totally unstable.
@@ 6.1.3 Comparison between different sGMRES variants with and without restarting
In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [Davis2011] (also tested in [guttel2024sketch]). The right-hand side vector is randomly generated according to the standard normal distribution. We run $\texttt{nrestarts}=5$ cycles of restarted sGMRES with the truncated Arnoldi process, and sketch-and-select Arnoldi, with $m=150$ iterations per cycle, totaling $750$ iterations. We also run restarted GMRES with the same parameters. Additionally, we perform sGMRES with truncated Arnoldi and sketch-and-select Arnoldi with $m=750$ iterations (without restarts). All implementations of sGMRES use a truncation parameter $t=3$.
@@
6.2 sGMRES with the adaptive restarting strategy
In the final experiment, we test the adaptive restarting strategy discussed in Section 3. In particular, we test
@@
  • restarted sGMRES: restart every $m$ iterations with fixed parameter $t=1$.
  • restarted sGMRES with adapt $t$: restart every $m$ iterations with adaptive parameter $t$ beginning with $t=1$ and $\texttt{tol}_{\tau}=\bm{u}$, i.e., Algorithm 4.
Figures 4 and 5 illustrate tests of three algorithms on matrices with different condition numbers $5.49\times 10^{3}$ and $9.64\times 10^{11}$.In each example, the right-hand side is generated randomly following the standard normal distribution, with $m$ values of $50$ and $100$, and $\texttt{nrestarts}=10^{3}/m$.
In Figure 4, both restarted GMRES and sGMRES achieve small backward errors, with restarted sGMRES using the adaptive restarting strategy accelerating convergence compared to a fixed $t$ approach by using a slightly larger $t$. As shown in the middle subfigure of Figure 4, it can be expected that the performance of restarted sGMRES, using an adaptive restarting strategy, tends towards the performance of restarted sGMRES.
Figure 5 clearly shows that, for $m=100$, our adaptive restarting strategy can enhance the poor backward error that arises from a small $t$ by adaptively increasing $t$. This effect is even more pronounced for $m=400$. Conversely, when $m=50$, the backward error is limited by the small $m$, not $t$, since the behavior of restarted GMRES mirrors that of restarted sGMRES with a small $t$, which does not offer any improvement for this scenario. In this case, $t$ is repeatedly doubled up to $m$ during the early stage, thus restarted sGMRES with the adaptive strategy performs similarly to standard GMRES. By comparing the experiments shown in Figures 4 and 5, we observe that restarted GMRES also requires a large $m$ in order to obtain a small backward error. According to Theorem 2, the reduced factor $\Lambda^{(i)}_{1}$ defined in (46) depends on $\bar{\tau}_{k^{(i)}}$, whose upper bound is determined by $\kappa(A)$. Therefore, for well-conditioned matrices such as fs_760_1, achieving $\Lambda^{(i)}_{1}\ll 1$ is easy, meaning that a relatively small $m$ typically suffices. In contrast, for ill-conditioned matrices, a significantly larger $m$ may be necessary.
Furthermore, in Figure 6, we combine the preconditioners generated by the MATLAB command ilu with the adaptive strategy to test sherman2, i.e., the same matrix tested in Figure 5. Due to the efficiency of the preconditioners, we can reduce the condition number of the preconditioned system and then employ much smaller $m$, that is, $m=5$, $10$. It can be seen that when preconditioning is used, restarted sGMRES with and without the adaptive strategy have similar behavior and require fewer orthogonalizations than standard GMRES.
In general, sGMRES with our adaptive restarting approach often performs comparably to standard GMRES regarding the accuracy, and has the potential to result in substantial time savings especially for the well-conditioned case.
Figure 4: Adaptive restarting experiment. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for fs_760_1, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
Figure 5: Adaptive restarting experiment. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
Figure 6: Adaptive restarting experiment with preconditioning. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.